OpenStudy (anonymous):

how to slove : q/5=41.

OpenStudy (anonymous):

multiply 5 on both sides

OpenStudy (anonymous):

but i thought , since the 5 is on the bottom you had to do somthing different...

OpenStudy (anonymous):

no that is correct if you take 205/5 you would get 41 which is 41*5 so it works

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

how about this problem: 50q-43=52q-81

OpenStudy (anonymous):

ok first thing you need to do is get your "q's" on one side and your other numbers on the other.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

write out what you got and i will check it

OpenStudy (anonymous):

19

OpenStudy (anonymous):

correct!1

OpenStudy (anonymous):

okay thanks ! i have more question, do you mind helping .. ?

OpenStudy (anonymous):

i would get in the practice of writing it out as q=19 it will save you problems in the long run

OpenStudy (anonymous):

multiply both sides with 5 q/5 * 5 = 41 * 5 q = 205

OpenStudy (anonymous):

whats your next problem?

OpenStudy (anonymous):

50q-43=52q-81 for this ur answer is : transpose 52q to LHS ( light hand side ) 50q-52q - 43 = -81 transpose -43 to right hand side -2q =- 81+43 -2q = -38 q = (-38)/(-2) q = 19 hope u got it

OpenStudy (anonymous):

ask ur next problem basket ball